For an AP the sum of the 4th and 5th terms is 32 and sum of 2nd and 3rd terms is 16. What is the common difference and first term?
Answer by chndnbl(31)

T4 + T5 = 32
a +3d + a+ 4d = 32
2a + 7d = 32......(1)
T2 + T3 = 16
a+ d + a+ 2d = 16
2a + 3d = 16......(2)
solving these equation by elimination method,
4d = 16
d =4